. a) How much more solar radiation is received on a surface when the sun angle is 70° compared to when the angle is 40°?

b) How much shorter is the sun’s path length through the atmosphere when the sun angle is 70° compared to when it is 40°?
1
Expert's answer
2020-10-15T03:13:46-0400

a)


"\frac{P_1}{P_2}=\frac{\sin{70}}{\sin{40}}=1.46"

b)


"\frac{L_1}{L_2}=\frac{\cos{70}}{\cos{40}}=0.45"